OpenStudy (bekkah323):

find the intercepts: 2y=6-3x

OpenStudy (rmrjr22):

put into y = mx+b

OpenStudy (rmrjr22):

then set y to 0 to find x

OpenStudy (bekkah323):

is it y=3/2x+3

OpenStudy (rmrjr22):

watch ur signs

OpenStudy (bekkah323):

y=-3/2x+3

OpenStudy (mertsj):

Find the x intercept by replacing y with 0 Find the y intercept by replacing x with 0

OpenStudy (rmrjr22):

now set y or x = to isolate the intercept

OpenStudy (rmrjr22):

= 0 **

OpenStudy (bekkah323):

x=2

OpenStudy (rmrjr22):

yes

OpenStudy (bekkah323):

i feel like i got this one wrong y=3/2

OpenStudy (rmrjr22):

just a lil

OpenStudy (rmrjr22):

y = m(0) + b

OpenStudy (rmrjr22):

y = b

OpenStudy (bekkah323):

ok, yah i see what i did wrong. thank you!!

OpenStudy (rmrjr22):

np